A flywheel rotates with a constant retardation due to breaking from t=0to t=10 seconds. it made 400 revolutions . At time t=8 sec , it's angular velocity was 45πrad/sec. Determine


1 . Value of constant ratardation.


2. Total time taken to come to rest.


3. Total revolution made till it comes to rest.


1
Expert's answer
2021-05-05T07:40:28-0400

"w = 2\pi*N/60"

"= 2\pi *400/60"

= 41.89 rad/sec

"= (45-41.89)\/2"

"1. = -1.556 rad\/sec^2"

"t =" 41.89/1.559

"2. time = 26.92 sec"

Number of revolution made before rest,